# Causal Inference for a Hidden Treatment

### v2

@@ 9 Discussion
In this paper, we considered the fundamental problem of causal inference with a hidden binary treatment, an all too common problem in epidemiology and the social sciences. We proposed a unified solution which delivers nonparametric identification of a rich class of causal parameters, by drawing on a strategy which carefully leverages a surrogate and a proxy variable of the hidden treatment to restore identification, while bypassing the need for validation data or restrictive functional form assumptions prominent in prior measurement error literature. An important contribution of the paper is to develop formal semiparametric theory which gives rise to influence function-based nonparametric estimators with important robustness and efficiency properties. The approach provides a generic roadmap for constructing novel semiparametric estimators in a large class of hidden treatment models. The paper also addresses the challenging issue of estimating high-dimensional nuisance functions which are defined in terms of the hidden treatment, therefore ruling out direct application of standard machine learning and nonparametric estimators. To resolve this difficulty, we proposed a semiparametric EM algorithm for estimating nuisance functions. We establish the semiparametric efficiency bound for discrete outcomes, and provide an approximation for continuous outcomes.

### v3

@@ 9 Discussion
    In this paper, we considered the fundamental problem of causal inference with a hidden binary treatment, an all too common problem in epidemiology and the social sciences. We proposed a unified solution which delivers nonparametric identification of a rich class of causal parameters, by drawing on a strategy which leverages a surrogate and a proxy variable of the hidden treatment to restore identification, while bypassing the need for validation data or restrictive functional form assumptions prominent in prior measurement error literature. An important contribution of the paper is to develop formal semiparametric theory which gives rise to influence function-based nonparametric estimators with important robustness and efficiency properties. The approach provides a generic roadmap for constructing novel semiparametric estimators in a large class of hidden treatment models. The paper also addresses the challenge of estimating high-dimensional nuisance functions defined in terms of the hidden treatment, therefore ruling out direct application of standard machine learning and nonparametric estimators. To resolve this difficulty, we proposed a semiparametric EM algorithm for estimating nuisance functions. We establish the semiparametric efficiency bound for discrete outcomes, and provide an approximation for continuous outcomes.